OpenStudy (anonymous):

How do you solve by substitution? X=2y+2 and 2x+3y=11

OpenStudy (anonymous):

You plug in what you have for x into the equation for instance in the equation 2x+3y=11 You'd plug in what's equal to X into the X part of that equation

OpenStudy (anonymous):

\[2(2y+2)+3y=11\]

OpenStudy (anonymous):

^^^shes right

OpenStudy (anonymous):

2(2y+2)+3y=11 4y+4+3y=11 7y=7 y=1